SFP Question # 4

 

Assets= 300% of total liabilities

Liabilities= 50% of Owner's Equity

Owner's Equity= 80% of total Assets

Difference between Owner's Equity and Liabilities = 2,000,000 pesos

 

How much is the total liabilities?


1
Expert's answer
2021-09-26T08:41:49-0400

Solution:

The accounting equation: Assets = Liabilities + Owners Equity

Let Owners Equity be X

Liabilities = 50"\%" of X = 0.5X

X – 0.5X = 2M

0.5X = 2M

X = 4,000,000 pesos

Owners Equity = 4,000,000 pesos

Liabilities = 50"\% \times" 4M = 2,000,000 pesos

Assets = 2M + 4M = 6,000,000 pesos